# Transformation of orientation and rotation angles of synchronous satellites: Application to the Galilean moons

### v1

@@ Abstract
The orientation and rotation of a synchronous satellite can be referred to both its Laplace plane and the ICRF equatorial plane, in terms of Euler angles or spin axis Cartesian coordinates and Earth equatorial coordinates, respectively. We computed second-order analytical expressions to make the transformation between the two systems and applied them to the Galilean satellites (Io, Europa, Ganymede, and Callisto). If one term of the spin axis Cartesian coordinates series is dominant, trigonometric series can be generated for the inertial and orbital obliquities, node longitude and offset with respect to the Cassini plane. Since the transformation does not require any fit of amplitudes and frequencies on numerical series, the physical meaning of the frequencies is preserved from the input series and the amplitudes can be directly related to the geophysical parameters of interest. We provide tables for the coordinates and angles’ series assuming that the satellites are entirely solid, and considering two different orbital theories. The possible amplitude ranges for the main terms are also examined in the case where a liquid layer is assumed in the interior model. We use our transformation method to propose an updated IAU WG solution which would result in an improvement with respect to zero obliquity models used so far. This method will also be useful for the interpretation of future Earth-based radar observations or JUICE data.
Icarus Keywords: Rotational dynamics; Satellites, dynamics; Jupiter, satellites
@@ 1 Introduction
The synchronous rotation of the Jupiter’s Galilean satellites, Io, Europa, Ganymede, and Callisto, was only confirmed with certainty after decades of observations of their surface features from Earth telescopes, see Camichel et al. [1943], Dollfus and Murray [1974].
@@
The Galilean satellites are assumed to be in the so-called Cassini state [Colombo, 1966, Peale, 1969], which means that their rotation axis follows the long-term precession of their orbit normal while their rotation about the spin axis is periodically modified by physical libration, due to the torque exerted on their triaxial shape by Jupiter. The derivation of control networks for the Galilean satellites from images taken by Voyager 1 and 2 and Galileo did not reveal any significant deviation between their spin and orbit axes [Davies and Katayama, 1981, Davies et al., 1998]. The obliquity of the satellites, which is the angular separation of the rotation axis and the orbit normal, is therefore expected to be small.
The obliquity and the libration amplitude of Ganymede will be measured with exquisite precision by JUICE’s 3GM experiment during the Ganymede orbital phase ($0.2$ and $0.4$ arcsec, respectively, see Cappuccio et al. [2020]), slightly better than the expected accuracy with the GALA instrument (between $0.5$ and $1.7$ arcsec, see Steinbrügge et al. [2019]). The obliquity of Europa cannot be measured by 3GM due to the limited data provided by the close flyby, whereas the obliquity of Callisto can be determined to within $0.06^{\circ}-0.3^{\circ}$ [Cappuccio et al., 2022].
@@
The Europa Clipper spacecraft’s radio science and camera instruments will measure the orientation of Europa spin axis [Mazarico et al., 2023, Turtle et al., 2024] with a precision of the order of $0.05^{\circ}$ and $0.001^{\circ}$, respectively. In parallel with space missions, the radar speckle tracking observations from Goldstone Solar System Radar and Green Bank Telescope between 2011 and 2023 provided estimates of the orientation of Europa and Ganymede’s rotation axis with an accuracy of $0.01^{\circ}$, Margot [2025]. Note that on Europa and Ganymede, at the equator, $1$ arcsec = $8$ m and $13$ m, while $0.0001$ deg = $3$ m and $5$ m, respectively.
To make the most of these measurements, the rotation models for the satellites must be accurate, both in their description of the link between observables and internal parameters and in the orbital model used. This precise modeling has been an ongoing effort within the rotation community for about 20 years (e.g., Henrard [2005], Bills [2005], Van Hoolst et al. [2008], Noyelles [2009], Rambaux et al. [2011], Baland et al. [2012], Van Hoolst et al. [2013], Baland et al. [2016], Coyette et al. [2025]) and has certainly not yet reached its conclusion.
@@
The orientation and rotation of a body can be described using at least two different sets of angles. The first set is the Euler angles with respect to a fixed plane with the inertial obliquity $\theta$, the node longitude $\psi$ and the prime meridian location $\phi$. This fixed plane can be the equatorial plane of the central planet at J2000 or, preferably, the Laplace plane. The second set of angles is the Earth equatorial coordinates with respect to the Earth equator at J2000 (= International Celestial Reference Frame ICRF equatorial plane): right ascension $\alpha$, declination $\delta$ and prime meridian location $W$ angle.
These coordinates are adopted by the IAU Working Group on Cartographic Coordinates and Rotational Elements (WGCCRE) to define a rotation model for the main solar system bodies, see Archinal et al. [2018].
@@
The goal of this study is to find analytical expressions to transform the orientation angles of the Galilean satellites between the Laplace plane and the ICRF equatorial plane, correct up to the second order in small parameters like the spin axis obliquity. This transformation can be applied if the orbital obliquity is assumed to be equal to 0 (orbit normal = spin axis) or finite and given by a theoretical model or by observations. It is therefore an improvement over the zero obliquity model, see for example Stark et al. [2018]. The method does not require a fit of amplitudes and frequencies on a numerical time series and preserves the physical meaning of the frequencies when the input quantities are given as series. Our transformation model can be used to define the future IAU WG solution for the rotation and orientation using right ascension, declination, and prime meridian location, as predicted by a theoretical model for the spin orientation. The link between the geophysical parameters of interest and the Earth equatorial coordinates will therefore be more direct than that obtained with a fit to a numerical series. This study is also useful for the interpretation of future Earth based observation or JUICE data, and can be applied to other satellites in synchronous rotation like Titan or the Uranian satellites.
Recently Yseboodt et al. [2023] described an analytical method that converts the Martian Euler angles into right ascension/declination and reciprocally, but this method can not be applied without any change for different reasons. First, the node longitude of the Galilean satellites shows a large precession rate (periods of 7, 30, 130 and 560 years) while the Martian method works better for periodic motion with a small precession rate (period of about 26 ky for Mars). Therefore the precession term can be considered as a first-order quantity for Mars but not for the Galilean satellites.
@@
Second, contrary to Mars obliquity $\varepsilon$ that is about $25^{\circ}$, the obliquity of the Galilean satellites is very small, probably much smaller than $1^{\circ}$. The denominators of the Mars transformation coefficients are usually proportional to $\sin\varepsilon$, which may give approximate first-order coefficients with large value. Lastly, the spin Cartesian coordinates $s_{x}$ and $s_{y}$ projected onto the Laplace plane are easily expressed as trigonometric series computed from the orbital forcing series and therefore are better suited as a starting point than the Euler angles $\theta$ and $\psi$ for a transformation to Earth equatorial coordinates.
@@ 2 Planes and angles
The transformation of a position $\vec{r}$ in the satellite body frame (BF) to its position in the Inertial Frame (IF) $\vec{R}$ is:
@@
with $\mathbf{M}$ the rotation matrix from the BF to the IF. The BF frame is attached to the satellite and centered at its center of mass. Its $x$ and $y$-axes define the satellite’s equator, and its $z$-axis defines its figure axis. Galilean satellites should experience polar motion due to the gravitational torque exerted by their parent planet [Coyette et al., 2025]. This is also the case for Titan, which in addition experiences polar motion due to the exchanges of angular momentum between a seasonally-varying atmosphere and the solid surface [Coyette et al., 2016, 2018]. However, for the sake of simplicity, we assume here that the figure axis coincides with the spin axis. Including polar motion in the transformation would simply require additional rotations, see for example Eqs. (1-3) of Yseboodt et al. [2023]. The IF is also centered at the center of mass, but its equatorial plane is the Earth mean equator at the J2000 reference epoch (or the equator of the International Celestial Reference Frame ICRF). The IFIF $X$-axis points towards the ascending node of the ecliptic on the Earth equator (i.e. towards the vernal point $IE$ in Fig. 1). The transformation matrix $\mathbf{M}$ can be written in two different ways:
@@
depending on whether Euler ($\theta,\psi,\phi$) or Earth equatorial ($\alpha_{S},\delta_{S},W$) angles are used for the spin position.
Figure 1: Planes and angles. Angles are not drawn to scale but are exaggerated for the purpose of illustration.
The orientation of the $z$-axis is either defined by the inertial obliquity $\theta$ and the node longitude $\psi$ with respect to the Laplace plane, or either by the right ascension $\alpha_{S}$ and declination $\delta_{S}$ with respect to the ICRF. The Laplace plane (LP) is the mean plane around which the instantaneous orbital plane of the satellite precesses, see Section 2.2. Note that here $\psi=\Omega_{IL\,LS}$ is the angle measured from the node of the Laplace plane on the ICRF, a convention that differs from that of Baland et al. [2012], where $\psi=\Omega_{QL\,LS}$ is measured from the node of the Laplace plane on the equator of Jupiter. The inertial obliquity $\theta$ of the spin axis is measured from the Laplace plane pole, and should not be confused with the orbital obliquity $\varepsilon$, measured from the orbit normal.
The prime meridian ($x$-axis of BF) is positioned by $\phi$ or $W$, depending on the angles’ set, see Fig. 1. The difference $W-\phi$ is defined as the angle $\beta$, which varies with time, see Section 4. The rotation angle $\phi$ of the satellite must include periodic variations, called librations, with respect to synchronous rotation. These variations are also included in $W$.
@@ 2.1 The orbital theory
The four Galilean satellites form a complex dynamical system. Their orbital variations, including precession, are mainly driven by the planet’s flattening on the one hand, and by mutual perturbations between the satellites and solar perturbations on the other hand, as well as by the Laplace resonance between Io, Europa and Ganymede. Current solutions for the ephemerides of the Galilean moons are mainly based on ground-based astrometric observations and space-based data from Voyager, Galileo, and Juno. For the numerical applications presented in this study (see Section 6), we choose as orbital theory for the Galilean satellites the numerical ephemerides JUP387 computed by R. Jacobson from JPL. It is available over the 160024001600-2400 time period. In Section 8.1, we quantify the changes in spin position obtained using the last version of NOE [Lainey et al., 2009] instead of JUP387.
Both JUP387 and the last version of NOE are represented by Chebyshev polynomials. However, a quasi-periodic decomposition of the orbital motion is particularly useful for computing the orientation/rotation of satellites. It is then possible to decompose the forcings into series and evaluate the rotational response term by term, retaining only the main terms of the solution. Using a Levenberg-Marquardt algorithm, we numerically fit the Laplace plane (see Section 2.2) on the 160022001600-2200 time period and the amplitudes, frequencies and phases of a periodic series
of at least $10$ terms for the projected coordinates of the orbit normal onto the Laplace plane ($n_{x}$, $n_{y}$, see Section 2.3). The frequencies correspond to integer combinations of the fundamental arguments of the dynamical system. We use the numerical values of Lainey et al. [2006] as a priori values for the fit. A linear trend is also adjusted for each angle.
The fit residuals are usually less than $0.0002^{\circ}$.
@@ 2.2 The reference plane
$1/\cos\delta_{LP}$ 2.3226 2.3236 2.3286 2.3508
$\tan\delta_{LP}\,/\cos\delta_{LP}$ 4.8686 4.8737 4.8971 5.0011
$1/2\,\tan\delta_{LP} -$1.0481 -1.0487 -1.0515 -1.0637
$\cos\delta_{LP}$ 0.4306 0.4304 0.4294 0.4254
$\sin\delta_{LP} -$0.9026 -0.9027 -0.9031 -0.9050
$1/4\sin 2\delta_{LP}$ 0.1943 0.1942 0.1939 0.1925
Table 1: Numerical values for the orientation angles of the Laplace plane for the four Galilean satellites and transformation coefficients, using the JUP387 ephemerides. Values in parentheses are from Archinal et al. [2018]. The Jupiter pole coordinates are $\alpha_{J}=268.0564^{\circ}$, $\delta_{J}=64.4953^{\circ}$.
As mentioned above, the rotation modeling community needs a fixed plane to serve as a reference plane for describing the orbital elements involved in the forcing torque and the resulting rotational response. The equator of Jupiter at epoch J2000 may be proposed as a fixed reference plane, as is usually the case to describe the orbital motion of the Galilean satellites. However, due to the influence of the Sun, Jupiter’s equator is not the best reference plane to describe the rotational response of the satellites (see e.g. Noyelles [2009]). A relevant choice is the plane perpendicular to the center of the orbital cone for a regular motion, or for non-regular motion, that keeps the orbit inclination variations as small as possible. This reference plane is known as the Laplace plane. Since there are different ways to compute the Laplace plane (depending, for example, on the choice of the ephemerides, the time interval, or the quantity to minimize), different numerical estimations could be obtained. Our numerical values of the Laplace plane position for the four Galilean satellites are given in Table 1, using the JUP387 ephemerides. The farther the satellite is from the planet, the farther the Laplace plane is from Jupiter’s equator.

### v2

@@ Abstract
The orientation and rotation of a synchronous satellite can be referred to both its Laplace plane and the ICRF equatorial plane, in terms of Euler angles or spin axis Cartesian coordinates and Earth equatorial coordinates, respectively. We computed second-order analytical expressions to make the transformation between the two systems and applied them to the Galilean satellites (Io, Europa, Ganymede, and Callisto). If one term of the spin axis Cartesian coordinates series is dominant, trigonometric series can be generated for the inertial and orbital obliquities, node longitude and offset with respect to the Cassini plane. Since the transformation avoids having to fit amplitudes and frequencies on the output numerical series, the physical meaning of the frequencies is preserved from the input series and the amplitudes can be directly related to the geophysical parameters of interest. We provide tables for the coordinates and angles’ series assuming that the satellites are entirely solid, and considering two different orbital theories. The possible amplitude ranges for the main terms are also examined in the case where a liquid layer is assumed in the interior model. We use our transformation method to propose an updated IAU WG solution which would result in an improvement with respect to zero obliquity models used so far. This method will also be useful for the interpretation of future Earth-based radar observations or Juice data.
Keywords: Rotational dynamics; Satellites, dynamics; Jupiter, satellites
Highlights
- We define the analytical transformation between Laplace plane and ICRF systems.
- We propose an updated orientation/rotation solution, based on a non-zero obliquity model.
- Second-order terms ensure transformation accuracy in the order of arcseconds.
- The transformation preserves the trigonometric form of the series.
@@ 1 Introduction
The synchronous rotation of the Jupiter’s Galilean satellites, Io, Europa, Ganymede, and Callisto, was only confirmed with certainty after decades of observations of their surface features from Earth telescopes, see Camichel et al. [1943], Dollfus and Murray [1974]. The Galilean satellites are assumed to be in the so-called Cassini state [Colombo, 1966, Peale, 1969], which means that their rotation axis follows the long-term precession of their orbit normal while their rotation about the spin axis is periodically modified (physical librations) due to the torque exerted on their triaxial shape by Jupiter. The derivation of control networks for the Galilean satellites from images taken by Voyager 1 and 2 and Galileo did not reveal any significant deviation between their spin and orbit axes [Davies and Katayama, 1981, Davies et al., 1998]. The obliquity of the satellites, which is the angular separation of the rotation axis and the orbit normal, is therefore expected to be small (at most a few degrees).
@@
The obliquity and the libration amplitude of Ganymede will be measured with exquisite precision by Juice’s 3GM experiment during the Ganymede orbital phase ($0.2$ and $0.4$ arcsec, respectively, see Cappuccio et al. 2020), slightly better than the expected accuracy with the GALA instrument (between $0.5$ and $1.7$ arcsec, see Steinbrügge et al. 2019). The obliquity of Europa cannot be measured by 3GM due to the limited data provided by the close flyby, whereas the obliquity of Callisto can be determined to within $0.06^{\circ}-0.3^{\circ}$ [Cappuccio et al., 2022]. The Europa Clipper spacecraft’s radio science and camera instruments will measure the orientation of Europa spin axis [Mazarico et al., 2023, Turtle et al., 2024, Steinbrügge et al., 2026] with a precision of the order of $0.05^{\circ}$ and $0.001^{\circ}$, respectively. In parallel with space missions, the radar speckle tracking observations from Goldstone Solar System Radar and Green Bank Telescope between 2011 and 2023 provided estimates of the orientation of Europa and Ganymede’s rotation axis with an accuracy of $0.01^{\circ}$, Margot [2025]. Note that on Europa and Ganymede, at the equator, $1$ arcsec = $8$ m and $13$ m, while $0.0001$ deg = $3$ m and $5$ m, respectively. To make the most of these measurements, the rotation models for the satellites must be accurate, both in their description of the link between observables and internal parameters and in the orbital model used. This precise modeling has been an ongoing effort for about 20 years (e.g., Henrard 2005, Bills 2005, Van Hoolst et al. 2008, Noyelles 2009, Rambaux et al. 2011, Baland et al. 2012, Van Hoolst et al. 2013, Baland et al. 2016, Coyette et al. 2026) and has certainly not yet reached its conclusion.
@@
The orientation and rotation of a body can be described using at least two different sets of angles. The first set is the Euler angles with respect to a fixed plane with the inertial obliquity $\theta$, the node longitude $\psi$ and the prime meridian location $\phi$. This fixed plane can be the equatorial plane of the central planet at J2000 or, preferably, the Laplace plane. The second set of angles is the Earth equatorial coordinates with respect to the Earth equator at J2000 (= International Celestial Reference Frame ICRF equatorial plane): right ascension $\alpha$, declination $\delta$ and prime meridian location $W$ angle. These coordinates are adopted by the IAU Working Group on Cartographic Coordinates and Rotational Elements (WGCCRE) to define a rotation model for the main solar system bodies, see Archinal et al. [2018].
@@
The goal of this study is to find analytical expressions to transform the orientation angles of the Galilean satellites between the Laplace plane and the ICRF equatorial plane, correct up to the second order in small parameters like the spin axis obliquity. This transformation can be applied if the orbital obliquity is assumed to be equal to 0 (orbit normal = spin axis) or finite and given by a theoretical model or by observations. It is therefore an improvement over the zero obliquity model, see for example Stark et al. [2018]. This method is particularly well suited to the transformation of periodic series, as it preserves the periodic form of the input series by avoiding a fit of amplitudes and frequencies on the numerical output time series. Our transformation model can be used to define the future IAU WG solution for the rotation and orientation using right ascension, declination, and prime meridian location, as predicted by a theoretical model for the spin orientation. The link between the geophysical parameters of interest and the Earth equatorial coordinates will therefore be more direct than that obtained with a fit to a numerical series. This study is also useful for the interpretation of future Earth based observation or Juice data, and can be applied to other satellites in synchronous rotation like Titan or the Uranian satellites.
@@
Recently Yseboodt et al. [2023] described an analytical method that converts the Martian Euler angles into right ascension/declination and reciprocally, but this method cannot be applied without any change for different reasons. First, the node longitude of the Galilean satellites shows a large precession rate (periods of -7, -30, -130 and -560 years) while the Martian method works better for periodic motion with a small precession rate (period of about 26 ky for Mars). Therefore the precession term can be considered as a first-order quantity for Mars but not for the Galilean satellites. Second, contrary to Mars obliquity $\varepsilon$ that is about $25^{\circ}$, the obliquity of the Galilean satellites is very small, probably much smaller than $1^{\circ}$. The denominators of the Mars transformation coefficients (see Eqs. (21) of Yseboodt et al. 2023) are usually proportional to $\sin\varepsilon$, which may give approximate first-order coefficients with large value. Lastly, the spin Cartesian coordinates $s_{x}$ and $s_{y}$ projected onto the Laplace plane are easily expressed as trigonometric series computed from the orbital forcing series and therefore are better suited as a starting point than the Euler angles $\theta$ and $\psi$ for a transformation to Earth equatorial coordinates.
@@ 2 Planes and angles
The transformation of a position vector $\vec{r}$ expressed in the satellite body frame (BF) into its representation $\vec{R}$ in the Inertial Frame (IF) is:
@@
with $\mathbf{M}$ the rotation matrix from the BF to the IF. The BF frame is attached to the satellite and centered at its center of mass. Its $x$ and $y$-axes define the satellite’s equator, and its $z$-axis defines its figure axis. The Galilean satellites should experience polar motion due to the gravitational torque exerted by their parent planet [Coyette et al., 2026]. This is also the case for Titan, which in addition experiences polar motion due to the exchanges of angular momentum between a seasonally-varying atmosphere and the solid surface [Coyette et al., 2016, 2018]. However, for the sake of simplicity, we assume here that the figure axis coincides with the spin axis. Including polar motion in the transformation would simply require additional rotations, see for example Eqs. (1-3) of Yseboodt et al. [2023]. The IF is also centered at the center of mass, but its equatorial plane is the Earth mean equator at the J2000 reference epoch (or the equator of the International Celestial Reference Frame ICRF). The IF $X$-axis points towards the ascending node of the ecliptic on the Earth equator (i.e. towards the vernal point $IE$ in Fig. 1). The transformation matrix $\mathbf{M}$ can be written in two different ways:
@@
depending on whether Euler ($\theta,\psi,\phi$) or Earth equatorial ($\alpha_{S},\delta_{S},W$) angles are used for the spin position. The rotation matrices are defined as
$R_{X}[w]=\left(\begin{array}[]{ccc}1&0&0\\ 0&\cos w&\sin w\\ 0&-\sin w&\cos w\\ \end{array}\right),\quad R_{Z}[w]=\left(\begin{array}[]{ccc}\cos w&\sin w&0\\ -\sin w&\cos w&0\\ 0&0&1\\ \end{array}\right).$ (3)
Figure 1: Planes and angles. Angles are not drawn to scale but are exaggerated for the purpose of illustration. The ICRF equator is denoted by $I$, the Jupiter equator by $Q$, the Laplace plane by $L$, the orbital plane by $O$, and the plane perpendicular to the spin axis by $S$. Nodes are the intersections between two planes and are labeled using the corresponding plane names. For example, QOQO denotes the node of the orbital plane on the Jupiter equator. The angle $\Omega_{N_{1}\,N_{2}}$ is measured from node $N_{1}$ to node $N_{2}$.
The orientation of the $z$-axis is either defined by the inertial obliquity $\theta$ and the node longitude $\psi$ with respect to the Laplace plane, or by the right ascension $\alpha_{S}$ and declination $\delta_{S}$ with respect to the ICRF. The Laplace plane (LP) is the mean plane around which the instantaneous orbital plane of the satellite precesses, see Section 2.2. Note that here $\psi=\Omega_{IL\,LS}$ is the angle measured from the node of the Laplace plane on the ICRF ($IL$), a convention that differs from that of Baland et al. [2012] and ending in $LS$, the node of the satellite equator on the Laplace plane, where $\psi=\Omega_{QL\,LS}$ is measured from the node of the Laplace plane on the equator of Jupiter ($QL$) up to $LS$. These two angles both belong to the Laplace plane. The inertial obliquity $\theta$ of the spin axis is measured from the Laplace plane pole, and should not be confused with the orbital obliquity $\varepsilon$, measured from the orbit normal.
The longitude systems of Europa, Ganymede and Callisto have been defined by reference to a crater by the IAU WGCCRE, whereas that of Io has been defined by the mean sub-Jovian direction (see Table 2 of Archinal et al. 2018). The prime meridian ($x$-axis of BF) is positioned by $\phi$ or $W$, depending on the angles’ set, see Fig. 1. The difference $W-\phi$ is defined as the angle $\beta$, which varies with time, see Section 4. The rotation angle $\phi$ of the satellite must include periodic variations, called librations, with respect to synchronous rotation. These variations are also included in $W$.
@@ 2.1 The orbital theory
The four Galilean satellites form a complex dynamical system. Their orbital variations, including precession, are mainly driven by the planet’s polar flattening on the one hand, and by mutual perturbations between the satellites and solar perturbations on the other hand, as well as by the Laplace resonance between Io, Europa and Ganymede. Current solutions for the ephemerides of the Galilean moons are mainly based on ground-based astrometric observations and space-based data from Voyager, Galileo, and Juno. For the numerical applications presented in this study (see Section 6), we choose as orbital theory for the Galilean satellites the numerical ephemerides JUP387 computed by R. Jacobson from JPL. They are available over the 160024001600-2400 time period. In Section 8.1, we quantify the changes in spin position obtained using the last version of NOE [Lainey et al., 2009] instead of JUP387.
Both JUP387 and the last version of NOE are represented by Chebyshev polynomials. However, a quasi-periodic decomposition of the orbital motion is particularly useful for computing the orientation/rotation of satellites. It is then possible to decompose the forcings into series and evaluate the rotational response term by term, retaining only the main terms of the solution. Using a Levenberg-Marquardt algorithm, we simultaneously fit the orientation of the Laplace plane (see Section 2.2) over the 160022001600-2200 time interval and a periodic series, with at least ten terms, describing the projected components of the orbit normal onto this plane ($n_{x}$, $n_{y}$, see Section 2.3). The fit provides the amplitudes, frequencies, and phases of the periodic terms. The frequencies correspond to integer combinations of the fundamental arguments of the dynamical system. The Laplace plane is given by the zero frequency term in our decomposition. We use the numerical values of Lainey et al. [2006] as a priori values for the fit. In order for each satellite to also follow the very slow (period of about 400 000400\,000 years, see Le Maistre et al. 2016) precession of Jupiter’s spin axis, a linear trend is also adjusted for each angle (see also Section 7). The trend may be difficult to separate from very long-period terms, such as Callisto’s 560-year orbital precession. However, for Callisto this periodic orbital perturbation has the dominant amplitude and is relatively large (0.25°, see Lainey et al. 2006), which helps to distinguish it from the secular trend. The fit residuals are usually less than $0.0002^{\circ}$, see Fig. 2.
Figure 2: Residuals of the orbit normal fit for the JUP and the NOE ephemerides for the four Galilean satellites, referred to their respective estimated Laplace plane.
@@ 2.2 The reference plane
$1/\cos\delta_{LP}$ $\mathcal{O}1$ 2.3226 2.3236 2.3286 2.3508
$\tan\delta_{LP}\,/\cos\delta_{LP}$ $\mathcal{O}2$ 4.8686 4.8737 4.8971 5.0011
$1/2\,\tan\delta_{LP}$ $\mathcal{O}2$ -1.0481 -1.0487 -1.0515 -1.0637
$\cos\delta_{LP}$ $\mathcal{O}1$ 0.4306 0.4304 0.4294 0.4254
$\sin\delta_{LP}$ $\mathcal{O}2$ -0.9026 -0.9027 -0.9031 -0.9050
$1/4\sin 2\delta_{LP}$ $\mathcal{O}2$ 0.1943 0.1942 0.1939 0.1925
$1/2\,\cos^{2}\delta_{LP}$ $\mathcal{O}2$ -0.0927 -0.0926 -0.0922 -0.0905
Table 1: Numerical values for the orientation angles of the Laplace plane for the four Galilean satellites and transformation coefficients, using the JUP387 ephemerides. The number of digits retained in this table is related to the order of magnitude of the residuals from the quasi-periodic decomposition of JUP387, see Fig. 2. $\mathcal{O}1$ denotes a first-order coefficient, while $\mathcal{O}2$ denotes a second-order coefficient (see Eqs. (9-10)). Values in parentheses are from Archinal et al. [2018]. The Jupiter pole coordinates are $\alpha_{J}=268.0564^{\circ}$, $\delta_{J}=64.4953^{\circ}$ (see ephemerides JUP387 computed by R. Jacobson from JPL).
As mentioned above, we need a fixed plane to serve as a reference plane for describing the orbital elements involved in the forcing torque and the resulting rotational response. The equator of Jupiter at epoch J2000 may be proposed as a fixed reference plane, as is usually the case to describe the orbital motion of the Galilean satellites. However, due to the influence of the Sun, Jupiter’s equator is not the best reference plane to describe the rotational response of the satellites (see e.g. Noyelles [2009]). A relevant choice is the plane perpendicular to the center of the orbital cone for a regular motion, or for non-regular motion, that keeps the orbit inclination variations as small as possible. This reference plane is known as the Laplace plane. Since there are different ways to compute the Laplace plane (depending, for example, on the choice of the ephemerides, the time interval, or the quantity to minimize), different numerical estimations could be obtained. Our numerical values of the Laplace plane position for the four Galilean satellites are given in Table 1, using the JUP387 ephemerides. The farther the satellite is from the planet, the farther the Laplace plane is from Jupiter’s equator.